# A Tutte-type canonical decomposition of 4-connected graphs

### v1

@@ 8.5. Angry block-bagels
Let $X$ be a graph. We call $X$ a double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ is a cycle. So $u$ and $v$ send edges to all vertices of $X-u-v$, and the edge $uv$ may or may not exist. The rim-length of $X$ then means the length of the cycle $X-u-v$. We call $X$ a double-wheel of triangles with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that $X-u-v$ has a cycle-decomposition $(O,\mathcal{G})$ with all adhesion-sets of size 1 and into triangles. The rim-length of $X$ means the length of the cycle $O$.
@@
Let $O$ denote the cycle $G-Z$. Let $e_{1},f_{1},e_{2},f_{2}$ be the first four edges on $O$. Let $(A_{i},B_{i})$ be the tetra-separation of $G$ with separator $\{e_{i},f_{i}\}\cup Z$. Then $(A_{1},B_{1})$ crosses $(A_{2},B_{2})$ with the four edges as links and with centre equal to $Z$. ∎
@@
Since $G$ is a double-wheel of triangles with rim-length $\geqslant 4$, the graph $G-Z$ has a cycle-decomposition $\mathcal{O}=(O,(G_{t})_{t\in O})$ with $|O|\geqslant 4$ such that all bags are triangles. Let $(A,B)$ and $(C,D)$ be crossing 2-separations of $O$. Let $\hat{A}:=Z\cup\bigcup_{t\in A}V(G_{t})$ and define $\hat{B},\hat{C},\hat{D}$ similarly. Then $(\hat{A},\hat{B})$ and $(\hat{C},\hat{D})$ are tetra-separations of $G$ that cross as desired. ∎
@@
(i) $\Rightarrow$ (ii). Assume that $G$ has two tetra-separations $(A,B)$ and $(C,D)$ that cross with all links of size one. Let $Z$ be the centre of $(A,B)$ and $(C,D)$, which is equal to the vertex-centre and has size two, by the (Crossing Lemma). (4.1). Let $\mathcal{O}$ be the block-bagel induced by $(A,B)$ and $(C,D)$. Then we have the (Block-bagel setting). (8.4). By the (Helpful Lemma). (8.6), we have $h(\mathcal{O})\geqslant 4$. If $\mathcal{O}$ has a good bag, then $G$ has a totally-nested tetra-separation by 8.27, contradicting that $G$ is 4-angry. Hence, every bag of $\mathcal{O}$ is bad.
(ii) $\Rightarrow$ (iii). Follows from Lemma 8.38.
(iii) $\Rightarrow$ (i). Follows from Lemma 8.39 and Lemma 8.40. ∎
@@ 9. Cycles of 2-connected torsos
One of the outcomes of Theorem 1 is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices. To investigate how this outcome arises, we assume that we are given two tetra-separations $(A,B)$ and $(C,D)$ of a 4-connected graph $G$ that cross with empty vertex-centre. In order to construct a cycle-decomposition for analysing the cyclic structure that is determined by $(A,B)$ and $(C,D)$, we first use the corner-diagram to divide $G$ into four smaller pieces (Section 9.1). We then apply the Tutte-decomposition to all four pieces and merge them to obtain a cycle-decomposition of $G$ (Section 9.2). In Section 9.3, we summarise assumptions shared by a fair share of the lemmas to come. Section 9.4 provides machinery that exploits the structure of the cycle-decomposition to find independent paths. We use this machinery in Section 9.5 to characterise which bags of the cycle-decomposition give rise to totally-nested tetra-separations (9.41). Section 9.6 shows that these totally-nested tetra-separations form a splitting star as in Theorem 1, and determines the structure of its torso (9.50). Finally, Section 9.7 characterises the special case where additionally all tetra-separations of the graph are crossed.
@@ 9.1. Bananas
Let $G$ be a 4-connected graph. Assume that two tetra-separations $(A,B)$ and $(C,D)$ of $G$ cross with all links of size two. Recall that the centre is empty by the (Crossing Lemma). (4.1). We obtain the graph ${\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{G^{*}}}$ from $G$ by adding edges, as follows. For each $L$-link $\{x,y\}$ and each adjacent corner $K=LS$ (where $L,S$ are sides), we join the unique vertices in $\hat{x}\cap S$ and $\hat{y}\cap S$ by an edge, which we call the $K$-edge for the $L$-link; see Figure 9.1. We will only use $G^{*}$ when $(A,B)$ and $(C,D)$ are clear from context. The expanded corner for $AC$ is obtained from the corner $(A\smallsetminus B)\cap(C\smallsetminus D)$ for $AC$ by adding the endvertices of all edges contained in the two adjacent links. The banana for $AC$, or $AC$-banana, is the subgraph of $G^{*}$ induced by the expanded corner for $AC$. The expanded corners and bananas are defined similarly for the other three corners.
@@
$\neg$ (i) $\Rightarrow$ $\neg$ (ii). Say $K$ is the corner for $AC$. Let $\beta$ denote the $AC$-banana. Let $e(A)$ denote the $AC$-edge for the $A$-link, and let $e(C)$ denote the $AC$-edge for the $C$-link. Assume that $e(A)=e(C)=:e$. We will show that $AC$ is potter. If both endvertices of $e$ lie in the corner for $AC$, then each link adjacent to $AC$ consists of a matching of size two with the endvertices of $e$ as one side of the matching; see Figure 9.2. As $G$ is 4-connected, the corner for $AC$ cannot contain any other vertices besides the endvertices of $e$. But then both endvertices of $e$ have degree $\leqslant 3$ in $G$, contradicting that $G$ is 4-connected. So we may assume that $e$ has an endvertex $v$ that lies in the $A$-link, say. Then the $C$-link contains an edge $f$ that ends in $v$. Hence $f$ dangles from the $A$-link through the $C$-link, which implies that the corner for $AC$ is potter by the (Crossing Lemma). (4.1).
$\neg$ (ii) $\Rightarrow$ $\neg$ (i) holds by definition. ∎
Figure 9.2. The situation in the proof of Lemma 9.2
@@
Figure 9.3. The construction of $(\hat{X},\hat{Y})$ in the proof of Lemma 9.3. The extensions of $X$ to $\hat{X}$ and of $Y$ to $\hat{Y}$ are indicated by the dashed lines.
@@ 9.2. Tutte-bananas and Tutte-bagels
Let $G$ be a graph. A 2-separation $(A,B)$ of $G$ is totally-nested if it is nested with every 2-separation of $G$. This is equivalent to saying that $(A,B)$ is half-connected and at least one of $G[A]$ and $G[B]$ is 2-connected [Tridecomp, Corollary 3.3.3]. Assume now that $G$ is 2-connected and $N$ denotes its set of totally-nested 2-separations. Then the tree-decomposition $\mathcal{T}=\mathcal{T}(N)$ defined by $N$ is the Tutte-decomposition of $G$. Every torso of $\mathcal{T}$ is 3-connected, a cycle or a $K_{2}$. Tutte introduced his decomposition in [TutteCon]. A modern proof is included in the appendix of [Tridecomp]. The Tutte-decomposition $\mathcal{T}=(T,\mathcal{V})$ has the following two properties:
@@
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. Let $K$ be one of the four corners. Recall that the $K$-banana is 2-connected by Lemma 9.3. The $K$-Tutte-banana is the Tutte-decomposition of the $K$-banana. The $K$-Tutte-path is the decomposition-tree of the $K$-Tutte-banana.
@@
(i). Let $u,v$ be the two vertices in $\Lambda$. Recall that $uv$ is an edge in $G^{*}$. So there is at least one bag $V_{t}\in\mathcal{V}$ that contains both $u$ and $v$. Let $V_{s}\in\mathcal{V}$ contain the two ends of the $AC$-edge for the $C$-link, and choose $V_{s}$ to minimise the distance from $s$ to $t$ in $T$. Recall that $T$ is a path by Lemma 9.6. If $t$ is not an end of $T$, there is $V_{r}\in\mathcal{V}$ with $t\in\mathring{r}Ts$ and $V_{r}\smallsetminus V_{t}$ non-empty, so every vertex in $V_{r}\smallsetminus V_{t}$ is separated from a vertex outside the $AC$-banana by the adhesion-set of $rt$ which has size two, contradicting 4-connectivity. Hence $t$ is an end of $T$. Similarly, either $u$ or $v$ avoids the adhesion-set of $V_{t}$, say $u$. If the torso of $V_{t}$ is a cycle, then $u\in S(C,D)$ violates the degree-condition. Thus the torso is 3-connected.
(ii). Since $\Lambda$ contains an edge $e$, the subgraph $H$ of $G^{*}$ induced by $\hat{\Lambda}$ is either a triangle or a 4-cycle, depending on whether the other element of $\Lambda$ besides $e$ is a vertex $v$ or an edge $f$, respectively. If $H$ is a triangle, then $v\in D\smallsetminus C$ has degree two in the $AC$-banana. Hence $H$ is a leaf-bag of $(T,\mathcal{V})$, and no other bag of $(T,\mathcal{V})$ includes $\Lambda$. ∎
The node $t$ of $T$ as in Lemma 9.7 is said to represent the $L$-link in $T$, and in the $K$-Tutte-banana.
@@
Without loss of generality, $K$ is the corner $AC$. Both the $A$-link and the $C$-link are represented by end-nodes $a$ and $c$ of $T$, respectively, by Lemma 9.7. Assume now that $a=c$. Then $T$ has no edge, since otherwise the adhesion-set of this edge would form a 2-separator of $G$, contradicting 4-connectivity. Hence $|T|=1$, and in particular the corner $AC$ is not potter.
We claim that both the $A$-link and the $C$-link consist of vertices only. Indeed, if the $A$-link contains an edge, then the $AC$-edge $e$ for the $A$-link is distinct from the $AC$-edge for the $C$-link by Lemma 9.2, so the ends of $e$ together form a 2-separator of the $AC$-banana. This 2-separator gives rise to a 2-separation of the $AC$-banana that cuts off the triangle-torso or 4-cycle-torso at $a$ by Lemma 9.7, and this 2-separation is totally-nested as its separator is spanned by the edge $e$. Hence the 2-separation corresponds to an edge of $T$, a contradiction. ∎
@@
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. For each corner $K$ let $\mathcal{T}^{K}=(T^{K},\mathcal{V}^{K})$ denote the $K$-Tutte-banana. Recall that $T^{K}$ is a path by Lemma 9.6.
@@
    we join the nodes $t$ and $t^{\prime}$ by an edge $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ in $O$ if the torsos associated with these nodes in their respective Tutte-bananas are 3-connected;
@@
    we otherwise identify the nodes $t$ and $t^{\prime}$ to a single node $)𝑜𝐿{\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{o(L)}}$ of $O$.
We say that $o(L)$ represents the $L$-link in $O$. For each node $t$ of $O$ that is not of the form $o(L)$, we define $G_{t}:=G[V^{K}_{t}]$ where $K$ is the unique corner with $t\in T^{K}$. For each node of $O$ of the form $o(L)$ for some link $L$, we consider the two nodes $t\in T^{K}$ and $t^{\prime}\in T^{K^{\prime}}$ that were identified to obtain $o(L)$, and we let $G_{o(L)}:=G[V^{K}_{t}]$ which is independent of $K$ by Lemma 9.7. We refer to the pair $(O,(G_{t}:t\in V(O))$ as the Tutte-bagel induced by $(A,B)$ and $(C,D)$. It remains to show that this really is a Tutte-bagel.
@@
Let $\mathcal{O}=(O,(G_{t}:t\in V(O))$ denote the Tutte-bagel induced by $(A,B)$ and $(C,D)$. First, we show that $\mathcal{O}$ is a cycle-decomposition. That $O$ is a cycle follows from Lemma 9.8.
(H1). Recall that there are no jumping edges and no diagonal edges by the (Crossing Lemma). (4.1). Every vertex and edge of $G$ is included in some expanded corner, and hence is included in the bag of some Tutte-banana. Hence every vertex and edge of $G$ appears in some $G_{t}$.
(H2). Let $v\in G$ be a vertex and consider $O_{v}:=O[\,\{t\in V(O):v\in G_{t}\}\,]$.
@@
Let $H_{t}$ be a torso of $\mathcal{O}$. If $t$ represents some link, then $H_{t}$ is a 4-cycle or a triangle. Otherwise $H_{t}$ stems from a Tutte-banana. Then $H_{t}$ is 3-connected, a 4-cycle or a triangle, by Lemma 9.6.

### v2

@@ 8.5. Angry block-bagels
Let $X$ be a graph. We call $X$ a double-wheel with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v is a cycle. So $u$ and $v$ send edges to all vertices of XuvX-u-v, and the edge $uv$ may or may not exist. The rim-length of $X$ then means the length of the cycle XuvX-u-v. We call $X$ a double-wheel of triangles with centre $\{u,v\}$ if $X$ is 4-connected and $u,v$ are distinct vertices of $X$ such that XuvX-u-v has a cycle-decomposition $(O,\mathcal{G})$ with all adhesion-sets of size 1 and into triangles. The rim-length of $X$ means the length of the cycle $O$.
@@
Let $O$ denote the cycle GZG-Z. Let $e_{1},f_{1},e_{2},f_{2}$ be the first four edges on $O$. Let $(A_{i},B_{i})$ be the tetra-separation of $G$ with separator $\{e_{i},f_{i}\}\cup Z$. Then $(A_{1},B_{1})$ crosses $(A_{2},B_{2})$ with the four edges as links and with centre equal to $Z$. ∎
@@
[…]
@@
(i) $\Rightarrow$ (ii). Assume that $G$ has two tetra-separations $(A,B)$ and $(C,D)$ that cross with all links of size one. Let $Z$ be the centre of $(A,B)$ and $(C,D)$, which is equal to the vertex-centre and has size two, by the (Crossing Lemma). (4.1). Let $\mathcal{O}$ be the block-bagel induced by $(A,B)$ and $(C,D)$. Then we have the (Block-bagel setting). (8.4). By the (Helpful Lemma). (8.6), we have $h(\mathcal{O})\geqslant 4$. If $\mathcal{O}$ has a good bag, then $G$ has a totally-nested tetra-separation by ˜8.27, contradicting that $G$ is 4-angry. Hence, every bag of $\mathcal{O}$ is bad.
(ii) $\Rightarrow$ (iii). Follows from Lemma˜8.38.
(iii) $\Rightarrow$ (i). Follows from Lemma˜8.39 and Lemma˜8.40. ∎
@@ 9. Cycles of 2-connected torsos
One of the outcomes of Theorem˜1 is a cycle of triangle-torsos and 3-connected torsos on $\leqslant 5$ vertices. To investigate how this outcome arises, we assume that we are given two tetra-separations $(A,B)$ and $(C,D)$ of a 4-connected graph $G$ that cross with empty vertex-centre. In order to construct a cycle-decomposition for analysing the cyclic structure that is determined by $(A,B)$ and $(C,D)$, we first use the corner-diagram to divide $G$ into four smaller pieces (Section˜9.1). We then apply the Tutte-decomposition to all four pieces and merge them to obtain a cycle-decomposition of $G$ (Section˜9.2). In Section˜9.3, we summarise assumptions shared by a fair share of the lemmas to come. Section˜9.4 provides machinery that exploits the structure of the cycle-decomposition to find independent paths. We use this machinery in Section˜9.5 to characterise which bags of the cycle-decomposition give rise to totally-nested tetra-separations (˜9.41). Section˜9.6 shows that these totally-nested tetra-separations form a splitting star as in Theorem˜1, and determines the structure of its torso (˜9.50). Finally, Section˜9.7 characterises the special case where additionally all tetra-separations of the graph are crossed.
@@ 9.1. Bananas
Let $G$ be a 4-connected graph. Assume that two tetra-separations $(A,B)$ and $(C,D)$ of $G$ cross with all links of size two. Recall that the centre is empty by the (Crossing Lemma). (4.1). We obtain the graph $\color[rgb]{0.6,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0.6,0,0}{G^{*}}$} from $G$ by adding edges, as follows. For each $L$-link $\{x,y\}$ and each adjacent corner $K=LS$ (where $L,S$ are sides), we join the unique vertices in $\hat{x}\cap S$ and $\hat{y}\cap S$ by an edge, which we call the $K$-edge for the $L$-link; see Figure˜9.1. We will only use $G^{*}$ when $(A,B)$ and $(C,D)$ are clear from context. The expanded corner for $AC$ is obtained from the corner $(A\smallsetminus B)\cap(C\smallsetminus D)$ for $AC$ by adding the endvertices of all edges contained in the two adjacent links. The banana for $AC$, or $AC$-banana, is the subgraph of $G^{*}$ induced by the expanded corner for $AC$. The expanded corners and bananas are defined similarly for the other three corners.
@@
$\neg$ ˜(i) $\Rightarrow$ $\neg$ ˜(ii). Say $K$ is the corner for $AC$. Let $\beta$ denote the $AC$-banana. Let $e(A)$ denote the $AC$-edge for the $A$-link, and let $e(C)$ denote the $AC$-edge for the $C$-link. Assume that $e(A)=e(C)=:e$. We will show that $AC$ is potter. If both endvertices of $e$ lie in the corner for $AC$, then each link adjacent to $AC$ consists of a matching of size two with the endvertices of $e$ as one side of the matching; see Figure˜9.2. As $G$ is 4-connected, the corner for $AC$ cannot contain any other vertices besides the endvertices of $e$. But then both endvertices of $e$ have degree $\leqslant 3$ in $G$, contradicting that $G$ is 4-connected. So we may assume that $e$ has an endvertex $v$ that lies in the $A$-link, say. Then the $C$-link contains an edge $f$ that ends in $v$. Hence $f$ dangles from the $A$-link through the $C$-link, which implies that the corner for $AC$ is potter by the (Crossing Lemma). (4.1).
$\neg$ ˜(ii) $\Rightarrow$ $\neg$ ˜(i) holds by definition. ∎
Figure 9.2. The situation in the proof of Lemma˜9.2
@@
[…]
@@ 9.2. Tutte-bananas and Tutte-bagels
Let $G$ be a graph. A 2-separation $(A,B)$ of $G$ is totally-nested if it is nested with every 2-separation of $G$. This is equivalent to saying that $(A,B)$ is half-connected and at least one of $G[A]$ and $G[B]$ is 2-connected [CK23, Corollary 3.3.3]. Assume now that $G$ is 2-connected and $N$ denotes its set of totally-nested 2-separations. Then the tree-decomposition $\mathcal{T}=\mathcal{T}(N)$ defined by $N$ is the Tutte-decomposition of $G$. Every torso of $\mathcal{T}$ is 3-connected, a cycle or a $K_{2}$. Tutte introduced his decomposition in [T66]. A modern proof is included in the appendix of [CK23]. The Tutte-decomposition $\mathcal{T}=(T,\mathcal{V})$ has the following two properties:
@@
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. Let $K$ be one of the four corners. Recall that the $K$-banana is 2-connected by Lemma˜9.3. The $K$-Tutte-banana is the Tutte-decomposition of the $K$-banana. The $K$-Tutte-path is the decomposition-tree of the $K$-Tutte-banana.
@@
˜(i). Let $u,v$ be the two vertices in $\Lambda$. Recall that $uv$ is an edge in $G^{*}$. So there is at least one bag $V_{t}\in\mathcal{V}$ that contains both $u$ and $v$. Let $V_{s}\in\mathcal{V}$ contain the two ends of the $AC$-edge for the $C$-link, and choose $V_{s}$ to minimise the distance from $s$ to $t$ in $T$. Recall that $T$ is a path by Lemma˜9.6. If $t$ is not an end of $T$, there is $V_{r}\in\mathcal{V}$ with $t\in\mathring{r}Ts$ and $V_{r}\smallsetminus V_{t}$ non-empty, so every vertex in $V_{r}\smallsetminus V_{t}$ is separated from a vertex outside the $AC$-banana by the adhesion-set of $rt$ which has size two, contradicting 4-connectivity. Hence $t$ is an end of $T$. Similarly, either $u$ or $v$ avoids the adhesion-set of $V_{t}$, say $u$. If the torso of $V_{t}$ is a cycle, then $u\in S(C,D)$ violates the degree-condition. Thus the torso is 3-connected.
˜(ii). Since $\Lambda$ contains an edge $e$, the subgraph $H$ of $G^{*}$ induced by $\hat{\Lambda}$ is either a triangle or a 4-cycle, depending on whether the other element of $\Lambda$ besides $e$ is a vertex $v$ or an edge $f$, respectively. If $H$ is a triangle, then $v\in D\smallsetminus C$ has degree two in the $AC$-banana. Hence $H$ is a leaf-bag of $(T,\mathcal{V})$, and no other bag of $(T,\mathcal{V})$ includes $\Lambda$. ∎
The node $t$ of $T$ as in Lemma˜9.7 is said to represent the $L$-link in $T$, and in the $K$-Tutte-banana.
@@
Without loss of generality, $K$ is the corner $AC$. Both the $A$-link and the $C$-link are represented by end-nodes $a$ and $c$ of $T$, respectively, by Lemma˜9.7. Assume now that $a=c$. Then $T$ has no edge, since otherwise the adhesion-set of this edge would form a 2-separator of $G$, contradicting 4-connectivity. Hence $|T|=1$, and in particular the corner $AC$ is not potter.
We claim that both the $A$-link and the $C$-link consist of vertices only. Indeed, if the $A$-link contains an edge, then the $AC$-edge $e$ for the $A$-link is distinct from the $AC$-edge for the $C$-link by Lemma˜9.2, so the ends of $e$ together form a 2-separator of the $AC$-banana. This 2-separator gives rise to a 2-separation of the $AC$-banana that cuts off the triangle-torso or 4-cycle-torso at $a$ by Lemma˜9.7, and this 2-separation is totally-nested as its separator is spanned by the edge $e$. Hence the 2-separation corresponds to an edge of $T$, a contradiction. ∎
@@
Let $G$ be a 4-connected graph. Let $(A,B)$ and $(C,D)$ be tetra-separations of $G$ that cross with all links of size two. For each corner $K$ let $\mathcal{T}^{K}=(T^{K},\mathcal{V}^{K})$ denote the $K$-Tutte-banana. Recall that $T^{K}$ is a path by Lemma˜9.6.
@@
[…]
@@
[…]
We say that $o(L)$ represents the $L$-link in $O$. For each node $t$ of $O$ that is not of the form $o(L)$, we define $G_{t}:=G[V^{K}_{t}]$ where $K$ is the unique corner with $t\in T^{K}$. For each node of $O$ of the form $o(L)$ for some link $L$, we consider the two nodes $t\in T^{K}$ and $t^{\prime}\in T^{K^{\prime}}$ that were identified to obtain $o(L)$, and we let $G_{o(L)}:=G[V^{K}_{t}]$ which is independent of $K$ by Lemma˜9.7. We refer to the pair $(O,(G_{t}:t\in V(O))$ as the Tutte-bagel induced by $(A,B)$ and $(C,D)$. It remains to show that this really is a Tutte-bagel.
@@
Let $\mathcal{O}=(O,(G_{t}:t\in V(O))$ denote the Tutte-bagel induced by $(A,B)$ and $(C,D)$. First, we show that $\mathcal{O}$ is a cycle-decomposition. That $O$ is a cycle follows from Lemma˜9.8.
˜(H1). Recall that there are no jumping edges and no diagonal edges by the (Crossing Lemma). (4.1). Every vertex and edge of $G$ is included in some expanded corner, and hence is included in the bag of some Tutte-banana. Hence every vertex and edge of $G$ appears in some $G_{t}$.
˜(H2). Let $v\in G$ be a vertex and consider $O_{v}:=O[\,\{t\in V(O):v\in G_{t}\}\,]$.
@@
Let $H_{t}$ be a torso of $\mathcal{O}$. If $t$ represents some link, then $H_{t}$ is a 4-cycle or a triangle. Otherwise $H_{t}$ stems from a Tutte-banana. Then $H_{t}$ is 3-connected, a 4-cycle or a triangle, by Lemma˜9.6.